four years hence a boy will be twice as old as he was 6 years ago. what is his present ages?
Answer by LinnW(1048)

Set x = current age
(x+4) is four years from now
(x-6) is six years ago
(x+4) = 2(x-6)
x+4 = 2x -12
add -4 to each side
x = 2x -16
add -2x to each side
-x = -16
x = 16
The boy is 16.